A rectangle's width is one-fifth its length, and its perimeter is 144m. Find the dimenision of the rectangle
Answer by checkley77(12844)

W=L/5
P=2L+2W
144=2L+2(L/5)
144=2L+2L/5 FIND THE LCD (5)
144=(5*2L+2L)/5
144=(10L+2L)/5 CROSS MULTIPLY
10L+2L=144*5
12L=720
L=720/12
L=60 ANS. FOR THE LENGTH.
W=60/5=12 ANS. FOR THE WIDTH.
PROOF:
144=2*60+2*12
144=120+24
144=144